OpenStudy (anonymous):

Solve the system. y = -1/3 x + 2 and x + 3y = 3 Answer (0, 1) (1, 0) (3, 1/3) (3/2, 0) no solution

OpenStudy (anonymous):

plug in y and see what you get

OpenStudy (anonymous):

I dont know how to do that....

OpenStudy (anonymous):

x+3(-1/3x+2)=3

OpenStudy (anonymous):

x-x+6=3 0=-3 Is that a true statement?

OpenStudy (anonymous):

I dont think so

OpenStudy (anonymous):

Well then, that means that there is no solution. Also if you plug in all the other points into the equation none of them satisfy the conditions.

OpenStudy (anonymous):

@charlieberzak does it make sense?

OpenStudy (anonymous):

yea thx, it does make sense, so the answer to this question is no solution?

OpenStudy (anonymous):

matrices BRAH

OpenStudy (anonymous):

Correct, and @chris00 I don't think he knows how to solve using matrices.. do you @charlieberzak ?

OpenStudy (anonymous):

no, not really :/

OpenStudy (anonymous):

oh. damn.

OpenStudy (jhannybean):

\[\large \color{red} {y=-\frac13 x +2}\]\[\large x+3\color{red}y=3\]\[\large x +3\left(-\frac13x +2\right)=3\]\[\large x -x +6 = 3\]\[\large 6 \neq 3\]No solution.

OpenStudy (anonymous):

Haha, nice job @Jhannybean as usual

OpenStudy (jhannybean):

Thank you!

OpenStudy (jhannybean):

does it make sense, @charlieberzak ?

OpenStudy (anonymous):

those two are parallel lines.

OpenStudy (anonymous):

and no worries @charlieberzak you'll understand matrices soon enough, and that's make life so much easier for you

OpenStudy (anonymous):

parallel and distinct lines!

OpenStudy (anonymous):

If you ever need help, just call Miss Jhanny :P

OpenStudy (anonymous):

looool

OpenStudy (anonymous):

and Chris

OpenStudy (anonymous):

and thisguy two

OpenStudy (anonymous):

and anyone else that's not me

OpenStudy (anonymous):

haha that made sense

OpenStudy (anonymous):

and john howard

OpenStudy (anonymous):

and btw long time no see @charlieberzak

OpenStudy (jhannybean):

You see how they are parallel lines?\[ \large\text{equation 1}: \ y= \color{blue}{-\frac13 }x+2\]\[\large \text{equation 2} : x+3y =3\]solve for y. \[\large 3y = -x +3 \]\[\large y=\color{blue}{ -\frac13} x + 3\]

OpenStudy (anonymous):

Pfft, her and her fancy equations and explanations xD

OpenStudy (jhannybean):

Are you @Luigi0210 ? O_o

OpenStudy (anonymous):

uhm, no

OpenStudy (anonymous):

parallel and DISTINCT lines.

OpenStudy (jhannybean):

Oh,you sound like him talking that way, lol.

OpenStudy (jhannybean):

distinct because they have different y-intercepts.

OpenStudy (anonymous):

haha chris we have already established that but thanks for the back up

OpenStudy (anonymous):

and you can't proof anything :D

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

who I am is of no importance

OpenStudy (anonymous):

for all i know u could be a fish

OpenStudy (anonymous):

exactly!